help with A=1/2h(B+b) (for b)
Answer by checkley71(8403)

A=1/2*H(B+b)
A=H(B+b)/2 NOW CROSS MULTIPLY
H(B+b)=2A
HB+Hb=2A
Hb=2A-HB
b=(2A-HB)/H